OpenStudy (anonymous):

can anybody solve this for me please: Use the basic identities to cahnge the expression to one involving only sines and cosines. Then simplify o a basic trigonometric function: (sinx)(tan x + cot x)

OpenStudy (anonymous):

\[\tan(x)+\cot(x)={\sin(x) \over \cos(x)}+{\cos(x) \over \sin(x) }={1 \over { \sin(x) \cos(x) } }\] Multiply the right side of the above by sin(x): \[{ \sin(x)\over{ \sin(x) \cos(x)} } = { 1 \over \cos(x) } = \sec(x)\]